An electric charge inside a sphere (r=0.81m) has a magnitude of +4.8x10^18C. Find the electric flux through this surface



1
Expert's answer
2022-03-08T08:56:41-0500

The electric flux through surface


"\Phi=\oint EdA=\frac{q}{\epsilon_0}""\Phi=\frac{4.8*10^{−18}}{8.85*10^{-12}}=5.42*10^{-7}\:\rm N\cdot m^2/C"